1: A horizontal, opaque flat surface at a steady-state temperature of 77C is exposed to an airflow having a freestream temperature of 27C with a velocity of 15 m/s, and a length of 1 m along the flow direction. The emissive power of the surface is 628 W/m2 , the irradiation from surroundingsis 1380 W/m2 , and the reflectivity is ?=0.40. (a) Determine the absorptivity (?) of the surface. Ans: 0.6 (b) Determine the net radiation heat transfer rate for this surface. Is this heat transfer to the surface or from the surface? Ans: -200 W/m2 (c) Determine the combined heat transfer rate for the surface (radiation+convection). Is this heat transfer to the surface or from the surface?
Answer #1

Heat transfer to the surface. flowing J opaque Sorsface L=E 1 = Ltet Vo e=0.40 2 = 0.6 Absorptivity (d, of the Surface Q = 0.Too v=15m/s © q cour a com a @radiation. Ts combined mode of Heat transfer hc (Ts-To) qrad = hy (Ts-To) (Ts-To) * A EAJ (Tst_Pr=1-85x105 x 1007 0.02551 =0.730 ju ul 43 1.95 x 0.029 (1467567.56) °.(0.730 hkle 2799.158 ke ,5 = 71.40 W/mtk lha =71.40 W/